# Enhancing the Harrow-Hassidim-Lloyd (HHL) algorithm in systems with large condition numbers

### v2

@@ I Introduction
In this work, we propose a new method to address the issue of handling $A$ matrices with large condition numbers, by introducing the post-selection-improved HHL (Psi-HHL) algorithm, which relies upon the intuition of extracting the correct signal by subtracting a wrong signal from a mixed signal (that contains both the correct and wrong signals). A signal, in the context of our work, refers to a scalar that is a feature of |$\ket{x}$. For the purposes of this work, we pick the overlap between the input and output states of HHL, which in the case of molecules corresponds to their correlation energies (up to some normalization constants), as our signal, although it is to be noted that Psi-HHL itself is applicable to features as general as transition matrix elements. The wrong signal is obtained by successfully post-selecting outcome bit ‘0’ instead of ‘1’ (we call this approach HHL1). The mixed signal is generated in an HHL quantum circuit by introducing an appropriately chosen $2\times 2$ unitary just before measurement of the HHL ancilla (we name this procedure HHL2). Figure 1 presents schematic diagrams outlining the workings of the HHL and the Psi-HHL algorithms.
@@
Figure 1: (a) Schematic of the HHL algorithm, where 1 is post-selected and for which $P(1)$ would be low for large $\mathcal{\kappa}$. |$\ket{b}$ is the input state register with $n_{b}$ qubits, while $n_{r}$ refers to the number of clock register qubits. The Hong-Ou-Mandel (HOM) module is a controlled-SWAP test without an additional ancilla qubit, and we use it to extract |$-\|\ket{b}_{\rm un}\|^{2}\ \|\ket{x}_{\rm un}\|\ \ |\langle b|x\rangle|$ | (interpreted as the correlation energy, $E_{\mathrm{c}}^{\mathrm{HHL}}$, when HHL is applied to molecular calculations), which is the feature of |$\ket{x}$ that we are interested in. Sub-figure (b) provides a block-diagram description of the HHL and the HOM modules. $QPE$ refers to the quantum phase estimation while $CR(\theta)$ refers to the controlled-rotation module that inverts the eigenvalues of $A$. The $CR(\theta)$ circuit is built out of uniformly controlled rotations involving a set of angles $\{\theta_{i}\}$, denoted compactly as just $\theta$ in the figure. The dashed cambered rectangle around the CNOT gate in the HOM circuit module, and which is accompanied by $(i)$, indicates that the gate is controlled on the ith qubit of the |$\ket{b}$ register. Sub-figure (c) consists of HHL1 in the top panel, where $P(0)$ is post-selected, and HHL2 in the bottom panel, where the traditional HHL module is appended with an additional mixing unitary, $R_{Y}(2\alpha)$, and where $P^{\prime}(1)$ is post-selected. Psi-HHL relies on subtracting the final outcomes of HHL1 and HHL2, and as the expression at the bottom of panel (c) indicates, we recover the correlation energy via the procedure to within a simple factor. In our notation, $E_{\mathrm{c}}^{\mathrm{HHL}}$ and $E_{\mathrm{c}}^{\mathrm{Psi-HHL}}$ are the correlation energies obtained from HHL and Psi-HHL respectively. Further, $a_{i},b_{i},c_{i},d_{i}\in\mathbb{R}$ are constants.
@@ II.1 The HHL algorithm
In the above expression, $\{\tilde{\lambda}_{i}\}$ refers to the estimates of the actual eigenvalues of $A$ ($1\geq\tilde{\lambda}_{i}\geq 1/\mathcal{\kappa}$) obtained via the QPE module of HHL. Typically, $C$ is chosen to be the minimum eigenvalue of $A$, but it is worth noting that incurring such a large classical overhead can always be circumvented via approaches such as AdaptHHL [2]. When one post-selects the outcome ‘1’ on the HHL ancilla qubit, one can think of it as choosing the ‘correct’ signal.
@@
For the sake of simplicity, we set this quantity hereafter as the feature of |$\ket{x}$ that is of interest to us, for the rest of this work, although Psi-HHL can be applied to more general features, such as transition matrix elements involving |$\ket{x}$. Therefore, even for non-quantum chemistry examples that we shall later consider, we will use the term ‘correlation energy’ to refer to the quantity from Eq. II.1. The number of trials needed to successfully post-select outcome `$1$’ is $\mathcal{O}(\mathcal{\kappa}^{2})$. Thus, in the event of large $\mathcal{\kappa}$, the success probability falls quickly, making it practically difficult to carry computations using large $\mathcal{\kappa}$ matrices. In this work, we demonstrate a new approach— ‘Psi-HHL’, towards handling large $\mathcal{\kappa}$ matrices by leveraging the probability of successfully post-selecting the ‘wrong’ outcome `$0$’, $P(0)$, in HHL.
@@ II.2 The Psi-HHL approach
In this section, we discuss the two essential ingredients— HHL1 and HHL2, which make up Psi-HHL as illustrated in Figure 1(c). The HHL1 module depicts the case where the ‘wrong’ signal is post-selected from the traditional HHL given in Figure 1(a) and Figure 1(b). The HHL2 entails traditional HHL module appended with a mixing $2\times 2$ unitary, given as $R_{Y}(2\alpha)$ and post-selecting 1 from the mixed signal. The role of this unitary is to mix the coefficients of $|0\rangle$ and $|1\rangle$ in the traditional HHL (see Eq. 1), thereby outputting a ‘mixed’ signal. The core idea behind Psi-HHL is to subtract the correlation energies associated with HHL1 and HHL2 thus leaving us with the correlation energy corresponding to the ‘correct’ signal. The gain in taking this route is the efficiency that we achieve in the number of shots expended. We summarize this in the following proposition:
@@ II.2.1 HHL1: Computing the wrong signal
The first HHL module, HHL1, serves to post-select the ‘wrong’ signal of the HHL ancilla qubit, that is, the state attached to |$\ket{0}$ in Eq.1. We re-write the state at the output of all the registers $|\psi_{\rm out}\rangle$ except the clock qubits register, as follows,
@@
Here, $|b\rangle_{\rm un}$ is the unnormalized input vector after amplitude encoding, and
@@ II.2.2 HHL2: Computing the mixed signal
The second HHL execution involves attaching the rotation gate $R_{Y}(2\alpha)$ on the HHL ancilla before the measurement as illustrated in the lower panel of Figure 1(c). This seemingly trivial addition to the algorithm introduces a new degree of freedom via the angle $\alpha$, enabling one to boost the probability of successful matrix inversion in the case of large $\mathcal{\kappa}$. We recall that $R_{Y}(2\alpha)$ is a rotation about the $Y$ axis on the Bloch sphere as given below
@@ II.3 Complexity aspects
In order to elaborate on the second part of our Proposition, we invoke the probability with which one post selects the desired outcome successfully in terms of $\kappa$.
@@
$P^{\prime}(1)$$\leq$ $\|[\sqrt{\mathbb{I}-C^{2}A^{-2}}\ sin(\alpha)$ (21)
@@
For the choice of $C=\gamma\lambda_{\min}$, where $\gamma\in(0,1)$, one can push the lower bound for the choice of given $\alpha$. This implies that fewer measurements are required to invert matrices with large condition numbers, as compared to the traditional HHL technique illustrated in Figure 1(a). We show this in Figure 2, where for a given $\alpha=60^{\circ}$, we obtain
@@ II.4 Singular matrices
In both the cases above $\vec{b}\in N(A)$, such that $N(A)$ is the null space of $A$ matrix. We execute HHL as well as Psi-HHL for these two cases by setting $C/\lambda_{i}$=00 whenever the eigenvalues of $A$ are zero, that is, $\lambda_{i}=0$. The traditional HHL algorithm would expend all the supplied shots without being able to post-select outcome $1$, since $P(1)=0$. The HHL1 of Psi-HHL successfully post-selects outcome 00 with probability $P(0)=1$. On the other hand, the HHL2 module of Psi-HHL post-selects outcome $1$ with probability $\sin^{2}(\alpha)$, see Eq. 10. The data points presented in Figure 3 demonstrate these probabilities for the two examples considered. Both these HHLs yield the solution $|x\rangle=|b\rangle$. The overlap $|\langle x|b\rangle|$ from the HOM module in HHL1 and HHL2 both yield a value of 1, with all of the additional $\sin({\alpha})$ factors cancelling out (see Figure 1). Subtracting the two overlap values results in 0, as expected. Thus, the Psi-HHL approach proves powerful when $P(0)\gg P(1)$. We now consider the case when $\vec{b}$ is an eigenvector of $A$. Here too, one could still get the same result via Psi-HHL by choosing the evolution time, $t$, as a multiple of $2\pi$, which could be incorrect. Therefore, in order to reliably calculate the outputs from both the cases using Psi-HHL, one chooses a non-trivial $t$ in the QPE module, such that it is not an integral multiple of $2\pi$, in order to avoid the possibility of obtaining $\lambda_{i}=0$ for both the cases. Choosing a non-trivial $t$ will resolve this issue, in the following way:
@@ III Results and discussions
  • Toy $A$ matrices of sizes $4\times 4$, $8\times 8$, $16\times 16$, $32\times 32$, and $64\times 64$: For practical reasons associated with computational resources, we do not choose matrices whose $\mathcal{\kappa}$ grows with the $A$ matrix size. Instead, we generate a sequence of $A$ matrices with systematically increasing condition numbers for every system size. For example, for $4\times 4$ matrix size, we begin with a sparse and diagonally dominant matrix (the latter is not a requirement but simplifies matters for our illustrations) such that its smallest diagonal entry is $2^{-(n_{r}-1)}$ and the largest 1. Recall that $n_{r}$ is the number of QPE clock register qubits. We generate the sequence by varying $n_{r}$ from 2 to 20. In such a sequence, $\mathcal{\kappa}$ grows exponentially with $n_{r}$, thus enabling us to reach the failure point of HHL rather rapidly and with existing computational resources. We also note that changing the position of the minimum diagonal element of $A$ along the diagonal does not affect our analysis.
@@
All of our simulation results are obtained by building on the Qiskit program [21].
@@
Figure 4: Figure showing a comparison of the performance of HHL and Psi-HHL approaches for toy matrices. Sub-figures whose top panels present the average probabilities of HHL and Psi-HHL with varying values of $\mathcal{\kappa}$ for different $A$ matrix sizes: $4\times 4$ (sub-figure (a)), $8\times 8$ (sub-figure (b)), $16\times 16$ (sub-figure (c)), $32\times 32$ (sub-figure (d)), and $64\times 64$ (sub-figure (e)). The bottom panel of each sub-figure represents the average Percentage Fraction Differences (PFD) of the overlap, $E_{c}$, calculated with HHL and Psi-HHL with respect to the classical value versus $\mathcal{\kappa}$. Each data point results from 10 repeated runs of $10^{6}$ shots each. Also, all the computations have been done with $A$ not diagonal and $\vec{b}$ in unequal superposition, as described in the main text. We also note here that the approximate equality is due to the fit coefficients in the figure being rounded off to two decimal places.
@@ III.1 Toy matrices: $4\times 4$ through $64\times 64$
and $\vec{b}=\begin{pmatrix}0.10\\ 0.01\\ 0.20\\ 1.00\end{pmatrix}$. The sparsity, $s$, of $A$ is defined as the number of non-zero entries in the row with the most number of non-zero elements and is set to 2. Our choices for the other matrices, $8\times 8$ through $64\times 64$ can be found in Figure A1 of the Appendix. We have set the following values for sparsity: 3 for $8\times 8$, 4 for $16\times 16$, 5 for $32\times 32$, and 7 for $64\times 64$ matrices. Throughout, we have ensured that the eigenvalues of the $A$ matrix lies between $1/\mathcal{\kappa}$ and 1. We provide more examples for the $4\times 4$ matrix size in Section A1 of the Appendix.
Figures 4(a) through 4(e) provides our findings for $4\times 4$ through the $64\times 64$ examples considered. Each data point in the sub-figures is the average result obtained over 10 repetitions, with each repetition involving 1 million shots. We verify that the observed trends do not change with 50 repetitions by picking the $4\times 4$ matrix as a representative example (Figures A2 and A3 of Supplemental Material). We pick $\alpha=60^{\circ}$ for the Psi-HHL calculations. The corresponding data is given in Table A1 of the Appendix. Our results obtained using other choices of angles, $70^{\circ}$ and $80^{\circ}$, are provided in Tables A2 and A3 respectively. We now return to Figure 4, where each sub-figure provides two sets of data: one which compares the probabilities, $P(0)$, $P(1)$, $P^{\prime}(0)$, and $P^{\prime}(1)$, obtained by executing HHL and Psi-HHL, and the other providing the PFDs for the two methods. We choose condition numbers all the way till about 1 million. Figures A2 and A3 of the Appendix shows these results for $\alpha$ of $60^{\circ}$, $70^{\circ}$, and $80^{\circ}$, with $10^{5}$ and $10^{6}$ shots respectively. The figures show that our data fits for HHL align with the expected behaviour of $P(1)\sim\mathcal{O}(\mathcal{\kappa}^{-2})$. We also find that $P^{\prime}(1)$ scales as $\sim\mathcal{O}(1)$ for large condition numbers (with the fit parameter $a=0.75$ matching with the Y-axis intercept for large condition number values). For completeness, we verify that the fit probabilities add to one (see Table A4 of the Appendix).
@@
We now comment on the PFDs. We immediately see that HHL fails at some value of $\mathcal{\kappa}$ for the chosen number of shots (indicated by ‘nan’ values in Tables A1, A2, and A3 of the Appendix). This is due to the fact that as $P(1)$ becomes very small, the fraction of the total number of shots that yield 1 upon measurement tend to very small values. Therefore, the number of shots available for the HOM module become too small to give any reasonable result beyond a certain value of $\mathcal{\kappa}$. The bottom panel of each sub-figure of Figure 4 shows that the error bars, in which the upper and lower points of the bars reflect the largest and the smallest PFD values obtained using HHL, increases as the condition number increases, till it reaches a critical value of $\mathcal{\kappa}$, beyond which HHL fails. In contrast, we see that for Psi-HHL, even at $\mathcal{\kappa}\ \approx$ 1 million, the length of the error bar is about 17 percent.
We now verify that HHL does not fail upon increasing the number of shots, but the number requires to be about 1 billion in order to get a comparable range of PFDs as Psi-HHL, while Psi-HHL only incurs 10 million shots, as Figure 5 shows. In particular, HHL gives a maximum PFD of 1.51 percent and a minimum of $-2.98$ percent with a billion shots, whereas Psi-HHL gives a maximum and minimum PFD of 2.33 and $-2.97$ percent respectively with only 10 million shots.
We finally comment on our results from Figure 4 on the larger size matrices ($8\times 8$ through $64\times 64$), where for Psi-HHL, we consider $\alpha=60^{\circ}$. Figures A4 through A7 provides the data for $70^{\circ}$ and $80^{\circ}$ cases, with Figure A4 also giving the information for $10^{5}$ shots. Tables A5 through A8 in the Appendix provide the accompanying data. We observe that broadly, the trends are similar to what we find for the smaller $4\times 4$ cases. Figure A8 of the Appendix summarizes the situations where HHL fails for all of the toy matrices considered, as a heat map. We find that for 1 million shots, HHL starts returning ‘nan’ values between $\mathcal{\kappa}$ of $2^{8}$ and $2^{10}$. In view of the computational cost involved, we did not go all the way up to a condition number of $2^{20}$ for matrices above $4\times 4$ size, and we stopped much earlier for the $64\times 64$ case.
Refer to caption Refer to caption Refer to caption
@@
Figure 6: Plots showing the increase in $\mathcal{\kappa}$ with $A_{mol}$ matrix size, denoted as $m$, for the case of 6 (sub-figure (a)), 8 (sub-figure (b)) and 10 (sub-figure (c)) fixed occupied orbitals in the Sapporo-DKH3-DZP-2012 basis set. The left-most legend on top of sub-figure (a) shows black circles whose areas increase proportionally to the number of virtual orbitals considered.

### v3

@@ I Introduction
In this work, we propose a new method to address the issue of handling $A$ matrices with large condition numbers, by introducing the post-selection-improved HHL (Psi-HHL) algorithm, which relies upon the intuition of extracting the correct signal by subtracting a wrong signal from a mixed signal (that contains both the correct and wrong signals). A signal, in the context of our work, refers to a scalar that is a feature of |$\ket{x}$. For the purposes of this work, we pick the overlap between the input and output states of HHL, which in the case of molecules corresponds to their correlation energies (up to some normalization constants), as our signal, although it is to be noted that Psi-HHL itself is applicable to features as general as transition matrix elements. The wrong signal is obtained by successfully post-selecting outcome bit ‘0’ instead of ‘1’ (we call this approach HHL1). The mixed signal is generated from an HHL quantum circuit by introducing an appropriately chosen $2\times 2$ unitary just before measurement of the HHL ancilla (we name this procedure HHL2). Figure 1 presents schematic diagrams outlining the workings of the HHL and the Psi-HHL algorithms.
@@
Figure 1: (a) Schematic of the HHL algorithm, where 1 is post-selected and for which $P(1)$ would be low for large $\mathcal{\kappa}$. |$\ket{b}$ is the input state register with $n_{b}$ qubits, while $n_{r}$ refers to the number of clock register qubits. The Hong-Ou-Mandel (HOM) module is a destructive version of the controlled-SWAP test without an additional ancilla qubit, and we use it to extract $|\langle b|x\rangle|$, using which we evaluate |$-\|\ket{b}_{\rm un}\|^{2}\ \|\ket{x}_{\rm un}\|\ \ |\langle b|x\rangle|$ | (interpreted as the correlation energy, $E_{\mathrm{c}}^{\mathrm{HHL}}$, when HHL is applied to molecular calculations), which is the feature of |$\ket{x}$ that we are interested in. The subscript ‘un’ stands for un-normalized. Sub-figure (b) provides a block-diagram description of the HHL and the HOM modules. $QPE$ refers to the quantum phase estimation while $CR(\theta)$ refers to the controlled-rotation module that inverts the eigenvalues of $A$. The $CR(\theta)$ circuit is built out of uniformly controlled rotations involving a set of angles $\{\theta_{i}\}$, denoted compactly as just $\theta$ in the figure. The dashed cambered rectangle around the CNOT gate in the HOM circuit module, and which is accompanied by $(i)$, indicates that the gate is controlled on the ith qubit of the |$\ket{b}$ register. Sub-figure (c) consists of HHL1 in the top panel, where $P(0)$ is post-selected, and HHL2 in the bottom panel, where the traditional HHL module is appended with an additional mixing unitary, $R_{Y}(2\alpha)$, and where $P^{\prime}(1)$ is post-selected. Psi-HHL relies on subtracting the final outcomes of HHL1 and HHL2, and as the expression at the bottom of panel (c) indicates, we recover the correlation energy via the procedure to within a simple factor. In our notation, $E_{\mathrm{c}}^{\mathrm{HHL}}$ and $E_{\mathrm{c}}^{\mathrm{Psi-HHL}}$ are the correlation energies obtained from HHL and Psi-HHL respectively. Further, $a,b,c,d\in\mathbb{R}$ are constants.
@@ II.1 The HHL algorithm
In the above expression, $\{\tilde{\lambda}_{i}\}$ refers to the estimates of the actual eigenvalues of $A$ ($1\geq\tilde{\lambda}_{i}\geq 1/\mathcal{\kappa}$) obtained via the QPE module of HHL. Typically, $C$ is chosen to be the minimum eigenvalue of $A$, but it is worth noting that incurring the large classical overhead associated with computing it can always be circumvented via approaches such as AdaptHHL [2]. When one post-selects the outcome ‘1’ on the HHL ancilla qubit, one can think of it as choosing the ‘correct’ signal.
@@
For the sake of simplicity, we set this quantity hereafter as the feature of |$\ket{x}$ that is of interest to us, for the rest of this work, although Psi-HHL can be applied to more general features, such as transition matrix elements involving |$\ket{x}$. Therefore, even for non-quantum chemistry examples that we shall later consider, we will use the term ‘correlation energy’ to refer to the quantity from Eq. II.1.
The number of trials needed to successfully post-select outcome `$1$’ is $\mathcal{O}(\mathcal{\kappa}^{2})$. Thus, in the event of large $\mathcal{\kappa}$, the success probability falls quickly, making it practically difficult to carry computations using large $\mathcal{\kappa}$ matrices. In this work, we demonstrate a new approach— ‘Psi-HHL’, towards handling large $\mathcal{\kappa}$ matrices by leveraging the probability of successfully post-selecting the ‘wrong’ outcome `$0$’, $P(0)$, in HHL.
@@ II.2 The Psi-HHL approach
In this section, we discuss the two essential ingredients— HHL1 and HHL2, which make up Psi-HHL as illustrated in Figure 1(c). The HHL1 module depicts the case where the ‘wrong’ signal is post-selected from the HHL circuit, as shown in Figure 1(b) (Figure 1(a) illustrates the common case of post-selecting `$1$’). The HHL2 entails the traditional HHL module appended with a mixing $2\times 2$ unitary ($R_{Y}(2\alpha)$) on the HHL ancillary qubit just before measurement, and post-selecting `$1$’ from the ‘mixed’ signal. The role of this unitary is to mix the coefficients of $|0\rangle$ and $|1\rangle$ in the traditional HHL (see Eq. 1), thereby outputting a ‘mixed’ signal. The core idea behind Psi-HHL is to subtract the correlation energies associated with HHL1 and HHL2 thus leaving us with the correlation energy corresponding to the ‘correct’ signal. The gain in taking this route is the efficiency that we achieve in the number of shots expended. We summarize this in the following proposition:
@@ II.2.1 HHL1: Computing the wrong signal
The first HHL module, HHL1, serves to post-select the ‘wrong’ signal of the HHL ancilla qubit, that is, the state attached to |$\ket{0}$ in Eq.1. We re-write the state, $|\psi_{\rm out}\rangle$, at the output of all the registers except the clock qubits register, as follows,
@@
In the expression given above, $|b\rangle_{\rm un}$ is the unnormalized input vector after amplitude encoding, and
@@ II.2.2 HHL2: Computing the mixed signal
The second HHL execution involves attaching the rotation gate $R_{Y}(2\alpha)$ on the HHL ancillary qubit before the measurement as illustrated in the lower panel of Figure 1(c). This seemingly trivial addition to the algorithm introduces a new degree of freedom via the angle $\alpha$, enabling one to boost the probability of successful matrix inversion in the case of large $\mathcal{\kappa}$. We recall that $R_{Y}(2\alpha)$ is a rotation about the $Y$ axis on the Bloch sphere as given below
@@ II.3 Complexity aspects
In order to elaborate on the second part of our Proposition, we invoke the probability with which one post-selects the desired outcome successfully in terms of $\kappa$.
@@
$P^{\prime}(1)$$\leq$ $\rm\|[\sqrt{\mathbb{I}-C^{2}A^{-2}}\ sin(\alpha)$ (21)
@@
For the choice of $C=\gamma\lambda_{\min}$, where $\gamma\in(0,1)$, one can push the lower bound for the choice of given $\alpha$. This implies that fewer measurements are required to invert matrices with large condition numbers, as compared to the traditional HHL technique illustrated in Figure 1(a). We obtain
@@
We show this in Figure 2, where we pick $\alpha=60^{\circ}$, the choice of $A$ matrix is $A=\begin{pmatrix}0.25&0.00&0.00&0.00\\ 0.00&0.75&0.00&0.00\\ 0.00&0.00&0.50&0.00\\ 0.00&0.00&0.00&1.00\end{pmatrix}$, and with six different choices of $\vec{b}$ (the eigenvector associated with the minimum eigenvalue of $A$, the eigenvector associated with the maximum eigenvalue of $A$, eigenvector associated with two other eigenvalues of $A$, equal superposition of all eigenstates of $A$, and an arbitrary superposition of eigenstates of $A$).
@@ II.4 Singular matrices
In both the cases above $\vec{b}\in N(A)$, such that $N(A)$ is the null space of $A$ matrix. We execute HHL as well as Psi-HHL for these two cases by setting $C/\lambda_{i}$=00 whenever the eigenvalues of $A$ are zero, that is, $\lambda_{i}=0$. The traditional HHL algorithm would expend all the supplied shots without being able to post-select outcome $1$, since $P(1)=0$. The HHL1 of Psi-HHL successfully post-selects outcome 00 with probability $P(0)=1$. On the other hand, the HHL2 module of Psi-HHL post-selects outcome $1$ with probability $\sin^{2}(\alpha)$ (see Eq. 10). The data points presented in Figure 3 demonstrate these probabilities for the two examples considered. Both these HHLs yield the solution $|x\rangle=|b\rangle$. The overlap $|\langle x|b\rangle|$ from the HOM module in HHL1 and HHL2 both yield a value of 1, with all of the additional $\sin({\alpha})$ factors cancelling out (see Figure 1). Subtracting the two overlap values results in 0, as expected. Thus, the Psi-HHL approach proves powerful when $P(0)\gg P(1)$. We now consider the case when $\vec{b}$ is an eigenvector of $A$. Here too, one could still get the same result via Psi-HHL by choosing the evolution time, $t$, as a multiple of $2\pi$, which could be incorrect. Therefore, in order to reliably calculate the outputs from both the cases using Psi-HHL, one chooses a non-trivial $t$ in the QPE module, such that it is not an integral multiple of $2\pi$, in order to avoid the possibility of obtaining $\lambda_{i}=0$ for both the cases. Choosing a non-trivial $t$ will resolve this issue, in the following way:
@@ III Results and discussions
  • Toy $A$ matrices of sizes $4\times 4$, $8\times 8$, $16\times 16$, $32\times 32$, and $64\times 64$: For practical reasons associated with (classical) computational resources, we do not choose matrices whose $\mathcal{\kappa}$ grows with the $A$ matrix size. Instead, we generate a sequence of $A$ matrices with systematically increasing condition numbers for every system size. For example, for $4\times 4$ matrix size, we begin with a sparse and diagonally dominant matrix (the latter is not a requirement but simplifies matters for our illustrations) such that its smallest diagonal entry is $2^{-(n_{r}-1)}$ and the largest 1. Recall that $n_{r}$ is the number of QPE clock register qubits. We generate the sequence by varying $n_{r}$ from 2 to 20. In such a sequence, $\mathcal{\kappa}$ grows exponentially with $n_{r}$, thus enabling us to reach the failure point of HHL rather rapidly and with existing computational resources. We also note that changing the position of the minimum diagonal element of $A$ along the diagonal does not affect our analysis.
@@
All of our simulations are developed and executed on the Qiskit software development kit [21].
@@
Figure 4: Figure showing a comparison of the performance of HHL and Psi-HHL approaches for toy matrices. The top panels in each sub-figure present the average probabilities of HHL and Psi-HHL with varying values of $\mathcal{\kappa}$ for different $A$ matrix sizes: $4\times 4$ (sub-figure (a)), $8\times 8$ (sub-figure (b)), $16\times 16$ (sub-figure (c)), $32\times 32$ (sub-figure (d)), and $64\times 64$ (sub-figure (e)). The bottom panel of each sub-figure represents the average Percentage Fraction Differences (PFD) of the overlap, $E_{c}$, calculated with HHL and Psi-HHL with respect to the classical value versus $\mathcal{\kappa}$. Each data point results from 10 repeated runs of $10^{6}$ shots. Also, all the computations have been carried out with $A$ not diagonal and $\vec{b}$ in unequal superposition, as described in the main text. We also note here that the approximate equality is due to the fit coefficients in the figure being rounded off to two decimal places. For the Psi-HHL results, we set $\alpha$ to be 60 degrees.
@@ III.1 Toy matrices: $4\times 4$ through $64\times 64$
and $\vec{b}=\begin{pmatrix}0.10\\ 0.01\\ 0.20\\ 1.00\end{pmatrix}$. The sparsity, $s$, of $A$ is defined as the number of non-zero entries in the row with the most number of non-zero elements, and is set to 2. Our choices for the other matrices, $8\times 8$ through $64\times 64$ can be found in Figure A1 of the Appendix. We have set the following values for sparsity: 3 for $8\times 8$, 4 for $16\times 16$, 5 for $32\times 32$, and 7 for $64\times 64$ matrices. Throughout, we have ensured that the eigenvalues of the $A$ matrix lies between $1/\mathcal{\kappa}$ and 1. We provide more examples for the $4\times 4$ matrix size in Section A1 of the Appendix.
Figures 4(a) through 4(e) provides our findings for $4\times 4$ through the $64\times 64$ examples considered. Each data point in the sub-figures is the average result obtained over 10 repetitions, with each repetition involving 1 million shots. For the Psi-HHL results, we set $\alpha$ to be 60 degrees. Each sub-figure provides two sets of data: one which compares the probabilities, $P(0)$, $P(1)$, $P^{\prime}(0)$, and $P^{\prime}(1)$, obtained by executing HHL and Psi-HHL, and the other providing the PFDs for the two methods. We choose condition numbers all the way till about 1 million. The figures show that our data fits for HHL align with the expected behaviour of $P(1)\propto\mathcal{\kappa}^{-2}$. We also find that $P^{\prime}(1)$ scales as $\propto 1$ for large condition numbers (with the fit parameter $a_{1}=0.75$ matching with the Y-axis intercept for large condition number values). We also consider two more cases, where $A$ is diagonal but $\vec{b}$ is in an equal superposition, and $A$ diagonal and $\vec{b}$ in an unequal superposition. Furthermore, we also consider the situations where $\alpha$ is set to 70 and 80 degrees. The data for all of these cases are presented in Figures A2 and A3 of the Appendix, accompanied by Tables A1 through A3. We also verify that the observed trends do not change with 50 repetitions, for which we pick the $4\times 4$ matrix (with $\alpha=60^{\circ}$ for the Psi-HHL calculations) from Figure 4 as a representative example (see Figures A4 and A5 of the Appendix). The corresponding data is given in Table A4 of the Appendix. Results obtained using other choices of angles, $70^{\circ}$ and $80^{\circ}$, are provided in Tables A5 and A6 respectively. For completeness, we also verify that the fit probabilities from Figure 4 add to one (see Table A7 of the Appendix).
@@
We now comment on the observed PFDs from Figure 4. We immediately see that HHL fails at some value of $\mathcal{\kappa}$ for the chosen number of shots (indicated by ‘nan’ values in Tables A1 through A6 of the Appendix). This is due to the fact that as $P(1)$ becomes very small, the fraction of the total number of shots that yield 1 upon measurement tend to very small values. Therefore, the number of shots available for the HOM module become too small to give any reasonable result beyond a certain value of $\mathcal{\kappa}$. Our data also show that the error bars, in which the upper and lower points of the bars reflect the largest and the smallest PFD values obtained using HHL, increases as the condition number increases, till it reaches a critical value of $\mathcal{\kappa}$, beyond which HHL fails. In contrast, we see that for Psi-HHL, even at $\mathcal{\kappa}\ \approx$ 1 million, the length of the error bar is about 17 percent.
We finally comment on our results from Figure 4 on the larger size matrices ($8\times 8$ through $64\times 64$), where for Psi-HHL, we consider $\alpha=60^{\circ}$. Figures A6 through A9 provides the data for $70^{\circ}$ and $80^{\circ}$ cases, with Figure A6 also giving the information for $10^{5}$ shots. Tables A8 through A11 in the Appendix provide the accompanying data. We observe that broadly, the trends are similar to what we find for the smaller $4\times 4$ cases. Figure A10 of the Appendix summarizes the situations where HHL fails for all of the toy matrices considered, as a heat map. We find that for 1 million shots, HHL starts returning ‘nan’ values between $\mathcal{\kappa}$ of $2^{8}$ and $2^{10}$. In view of the computational cost involved, we did not go all the way up to a condition number of $2^{20}$ for matrices above $4\times 4$ size, and we stopped much earlier for the $64\times 64$ case.
We now turn our attention to Figure 5, where we see that HHL does not fail upon increasing the number of shots, but the number requires to be about 1 billion in order to get a comparable range of PFDs as Psi-HHL, while Psi-HHL only incurs 10 million shots. In particular, HHL gives a maximum PFD of 1.51 percent and a minimum of $-2.98$ percent with a billion shots, whereas Psi-HHL gives a maximum and minimum PFD of 2.33 and $-2.97$ percent respectively with only 10 million shots.
@@
Figure 6: Plots showing the increase in $\mathcal{\kappa}$ with $A_{mol}$ matrix size, denoted as $m$, for the case of 6 (sub-figure (a)), 8 (sub-figure (b)) and 10 (sub-figure (c)) fixed occupied orbitals in the Sapporo-DKH3-DZP-2012 basis set. The legend to the left at the top of the figure shows solid circles whose areas increase proportionally to the number of virtual orbitals considered.